11. Factorise : 27r + y3 + z8 - 9xyz.

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
SoL Consider ?:1.r3 + y3 + ii - 9xyz = (Sx)S + C y'jJ + )3 - 3<3xX yXz) [Using identity:.r3 + y3 + ii - 3%)'% = (.r + y + z)(r + y2 + z2 -%)' -yz - u)] POLYNOMALS 43 = (8% + Y + z) 1<8%>2 + y2 + z2 - (8%)y -yz - (8%)z} = (8% + y + z) <9%2 + y2 + z2 - 3x)i -yz - Su).